7x-20= 6y
-4x-2y= -44
solve the system

Answer by ilana(307)

First, we can rewrite -4x-2y=-44 as 2x+y=22 by dividing by -2. So y=22-2x. Substituting this y into 7x-20=6y gives 7x-20=6(22-2x). Distributing,
7x-20=132-12x. So 19x=152, so x=152/19=8.
Plugging this into y=22-2x gives y=6.